# Non-Hermitian entanglement dip from scaling-induced exceptional criticality

### v1

@@ I.2 The new approach: Scale-dependent energies and GBZ
    This approximate expression for the GBZ is dependent on the system size $L$ in the form of the exponent $1/(L+1)$, which gives weak dependence on the inter-chain coupling $\delta$ as well as a constant factor containing the other system parameters. While it is not the only possible approximation, it is relatively compact and agrees well with the numerical GBZ values [Fig. S1(b)] obtained by substituting the exact OBC spectrum $E_{0}$ of our 4-component physical system into $\text{det}(\mathcal{H}(z)-E_{0}\mathbb{I})=0$, and solving for $z$. Also, it predicts a spectrum $E_{\text{GBZ}}=\pm\sqrt{(t_{L}+1/z_{1})(t_{R}+z_{1})}$ that agrees very well with that numerically-obtained exact OBC spectrum $E_{0}$, as shown in Fig. S1(c-f). Regardless of the exact choice of approximation, the $\delta^{1/(L+1)}$ dependence holds universally, and suggests a slow but $\delta$-dependent convergence of the GBZ onto the unit circle in the $L\rightarrow\infty$ limit. In Fig. S1(c’-f’), the $|z_{1}|^{x},|z_{2}|^{x}$ spatial decay profiles predicted by the GBZ (brown) near $E_{0}=0$ also exhibit excellent agreement with the upper and lower envelope limits of the numerical eigenstate $\Psi$ closest to $E_{0}=0$ in Fig. S1(c-f).
In all, our approximate size-dependent GBZ solution for our coupled HN model system is given by
@@
is the transition length in which the energy spectrum or skin depth $-\log|z|$ switches from being size-independent to size-dependent [Fig. S1(b)]. The value of $L^{\prime}$ is determined by setting both expressions for the GBZ to be equal i.e. $|z_{1}|=\sqrt{t_{R}/t_{L}}=\exp\left(-{\alpha}/({L^{\prime}+1})\right)$. The very good agreement between our analytic GBZ solution (Eq. S44) and exact numerics is presented in Fig. S1 for the skin depth $-\log|z|$, spectrum $E$ and eigenfunction decay.
@@
Figure S1: Scaling-dependent GBZ, spectrum and eigenfunctions of the 4-component model $\mathcal{H}$ [Eq. S1]. (a) The two chain 4-component model [Eq. S1] incorporates equal but opposite hopping asymmetries $t_{L}$ and $t_{R}$ in both chains I and II. Even a small inter-chain coupling $\delta$ can lead to significant scaling dependence when $t_{L}\neq t_{R}$. (b) Agreement of the scale-dependent skin depth $\log|z|$ obtained via two different ways: numerically, by solving $\det(\mathcal{H}(z)-E_{0})=0$ from the numerical eigenspectrum $E_{0}$ and choosing slowest decaying $z$ solution (blue crosses); analytically, via the scale-dependent generalized Brillouin zone (GBZ) we derived [Eq. S44] (red dots) . $L^{\prime}$ [Eq. S45] represents the system size at which the skin depth (GBZ) transitions from being scale-independent to scale-dependent. (c-f, c’-f’) Energy spectra at various system sizes $L$ and illustrative eigenfunctions at energies $E_{0}$ (green diamonds in (c-f)) of smallest absolute value. Excellent agreement is observed between the numerical (blue) and analytically-derived GBZ (red) spectra [Eq. S44] and eigenfunctions [Eq. S22,Eq. S23]. The lower bound for the eigenfunctions (lower red curves) results from the superposition of $z_{1}$ and $z_{2}$ solutions in Eq. S23, and is generally non-monotonic except at the critical scale $L\approx L_{c}$ [Eq. LABEL:Eqcritical1], where it exhibits a perfectly exponential profile since both $z$ solutions coalesce. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$, and $\delta=1.6\times 10^{-3}$ for all panels.
@@
In a nutshell, we have shown that as the system size $L$ increases, the GBZ and spectrum always switches from being size-independent to size-dependent as $L$ crosses $L^{\prime}$. This is given by our approximate GBZ expression Eq. S44 which nevertheless exhibits excellent agreement with numerical results from exact diagonalization. Our approach was based on neglecting the effect of weak coupling terms in the bulk equations but retaining their influence in enforcing the boundary conditions, which allowed for analytic headway in writing down the energy expression from the characteristic equation of a single chain $\det(H(z)-E)=0$. This is justified because while extremely weak couplings should not significantly affect the bulk, they would still nontrivially interplay with the boundary conditions due to the high sensitivity of non-Hermitian systems to spatial inhomogeneities. To conclude, with nonzero coupling between chains with antagonistic NHSE, the GBZ is found to be no longer given by the usual $|z_{1}|=|z_{2}|$ condition, but instead takes on a $L$-dependent behavior that crucially involves the coupling strength $\delta$.
@@ I.3 Effective size-dependent two-component model
As such, working in our effect size-dependent GBZ $z$, we can accurately study $\mathcal{H}$ (Eq. S1) as two uncoupled copies of 2-component chains $H$ i.e. $\mathcal{H}=H(K)\oplus H^{T}(-K)$, where $K=-i\log z$. The effects of the coupling and NHSE are implicit in the imaginary part of $K$. From Eq. S2,
@@
from Eq. S44, where the real part $k=\text{Re}K=2m\pi/(L+1)$ with $m=1,2,...,L$, and the imaginary part $\text{Im}K=\displaystyle\frac{\alpha}{L+1}$ is controlled by
@@
from Eq. S44 and S45. From direct diagonalization of $H(K)$, the eigenenergies are given by
@@
which has a similar form as Eq. S3; but when evaluated on the GBZ (Eq. S44), it should reproduce the OBC spectrum of the entire original 4-component model $\mathcal{H}$ (Eq. S1). Note that we have used the notation $\mathcal{E}$ to refer to the eigenenergies predicted by the GBZ, which are supposed to closely approximate the true eigenenergies $E$.
@@
with $\alpha=-\log\left|\delta\times\displaystyle\frac{t_{R}(t_{R}-t_{L})}{t_{L}(t_{R}-1)}\right|$. Here, $L_{c}$ represents the critical length in which an EP appears, and should not be confused with $L^{\prime}=-\alpha/\log\sqrt{t_{R}/t_{L}}-1$ (Eq. S45), which is the length above which the GBZ becomes size dependent. Evidently, such a scaling-induced EP exists only if $L_{c}>L^{\prime}$.
@@
Figure S2: The critical system size $L_{c}$ dependence on $\delta$ for the coupled 4-component model [Eq. S1] and its effective two-component approximation [Eq. S46]. The spectral gap of our model closes at an EP at the critical system size $L_{c}$, which is analytically predicted to depend logarithmically on the small inter-chain coupling parameter $\delta$. Excellent agreement is observed between the numerically determined integer $L_{c}$ (black dots) and its analytical prediction [Eq. LABEL:Eqcritical1]. Since $\delta$ is built into the GBZ for the two-component, single-chain approximation [Eq. S46], its excellent numerical agreement is testimony to the validity of our scale-dependent GBZ approach.
@@ I.3.1 Critical behavior near the scaling-induced EP
To prepare for the study of the entanglement properties near the scaling-induced EPs described above, we describe how the Hamiltonian behaves in their neighborhood. We first expand the size-dependent quasi-momentum $K$, both its real and imaginary parts, as $\Delta_{K}=K-K_{c}$ where $K=k+i\alpha/(L+1)$ and $K_{c}=\pi+i\alpha/(L_{c}+1)$ [See Eqs. S44 and LABEL:Eqcritical1], such that $\Delta_{K}=\text{Re}\Delta_{K}+i\text{Im}\Delta_{K}$ is given by
@@
With these, the Hamiltonian (Eq. S46) can be expanded near $K=K_{c}$ as
$\begin{split}H(K)&\xrightarrow{\Delta_{K}=K-K_{c}}\begin{pmatrix}0&t_{L}+\text{e}^{-iK_{c}}\text{e}^{-i\Delta_{K}}\\ t_{R}+\text{e}^{iK_{c}}\text{e}^{i\Delta_{K}}&0\end{pmatrix}\\ &\xrightarrow[t_{R}+\exp(iK_{c})=0]{\text{e}^{i\Delta_{K}}\approx 1+i\Delta_{K}}\begin{pmatrix}0&t_{L}-1/t_{R}\\ -it_{R}\Delta_{K}&0\end{pmatrix}\ .\\ \end{split}$ (S52)
This yields the eigenenergies of $H(K_{c}+\Delta_{K})$ to be ±=±$\pm\mathcal{E}=\pm\sqrt{i(1-t_{L}t_{R})\Delta_{K}}$.
@@ I.3.2 Critical properties of the occupied band projector $P_{K}$
The geometric defectiveness of an EP is reflected in the divergences of the occupied band projector $P$, since the latter will not be well-defined with the occupied and valence bands merge into one. For a 2-component model, $P=\sum_{k}P_{k}|k\rangle\langle k|$ is defined (in the biorthogonal basis) as
@@
where $\mathcal{E}(K)$ is the eigenvalue of the Hamiltonian $H(K)$ (Eq. S46). Note that we have analytically continued $P_{k}\rightarrow P_{K}$ into the GBZ. Ordinarily, away from gap closure points, $P(K)$ projects onto a well-defined occupied band since $\mathcal{E}(K)\neq 0$. But at an EP ($K=K_{c}$)where $\Delta_{K}=0$, $\mathcal{E}(K)=\sqrt{i(1-t_{L}t_{R})\Delta_{K}}$ vanishes, rendering $P_{K}$ singular.
Close to an EP, where $\Delta_{K}$ is very small,
@@
Figure S3: Potentially divergent matrix element $P_{K}^{+-}$ [Eq. S54] of our truncated projection operator $P_{K}$ [Eq.S53] in even vs. odd $L$ systems. (a) Complex values of the off-diagonal element $P_{K(k)}^{+-}$, which traces a different loop for each system size $L$ as $k$ cycles over a period $[0,2\pi)$ with discrete points $(2m-1)\pi/L$, $m=1,2,...,L$ (drawn as solid dots). The $k=\pi$ point exist only for odd $L$, and causes $P_{K(k)}^{+-}$ to diverge to infinity as $L\rightarrow L_{c}=-\frac{\alpha}{\log t_{R}}-1=42.79$ [Eq. LABEL:Eqcritical1], $\alpha=-\log\left|\delta\times\displaystyle\frac{t_{R}(t_{R}-t_{L})}{t_{L}(t_{R}-1)}\right|$. (b, c) The values of $|P_{K(k)}^{+-}|$, Re$(P_{K(k)}^{+-})$ and Im$(P_{K(k)}^{+-})$ as a function of system size $L$, plotted as separate branches for each $k$ value for (b) even and (c) odd $L$. While there exists for even $L$ a $k=\pi-\pi/(L+1)$ branch that departs from the other branches, for odd $L$, the $k=\pi$ branch exhibits a sharp kink for $L=43\approx L_{c}=42.79$. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3},\delta=1.6\times 10^{-3}$.
@@ II Free-fermion entanglement entropy dip due to scaling-induced EPs
where ¯$\bar{\mathcal{R}}_{[x_{L},x_{R}]}$ is the projector onto sites outside of $[x_{L},x_{R}]$ (note that the index $n$ was expressed as indices $(k,\mu)$ in the main text, such as to emphasize the momentum $k$). For two-component models, the above general expression for $P$ reduces to Eq. S53. The eigenvalues of the projection operators can be interpreted as occupancy probabilities. For $P$, they are limited to either 1 and 0. However, when the system is divided in to the $[x_{L},x_{R}]$ and its complement, the truncated occupied matrix $\bar{P}$ reveals the entanglement between these two regions, resulting in occupancy probabilities (eigenvalues) $p$ away from $1$ and 00. The free-fermion entanglement entropy (EE) $S$ quantifies this entanglement, and is given by
@@
II.1 Entanglement entropy scaling and dip in the effective 2-component model Eq. S46
@@
Because of the peculiar scaling properties of our GBZ of our effective 2-band model Eq. S46, $\bar{P}$ behaves very differently for odd and even $L$ [See Fig. S4]. For odd $L$, we have shown that the discretized momenta in the GBZ contain one point where $ReK=\pi$, i.e. min$|\text{Re}\Delta_{K}|$ from the EP is zero, where $\Delta_{K}=i\left(\frac{\alpha}{L+1}-\frac{\alpha}{L_{c}+1}\right)$. In this case, $P^{+-}_{K}$ [Eq. S55] is divergent as $|L-L_{c}|^{-1/2}$ near the EP. However, for even $L$, the divergence is only asymptotic with $L^{1/2}$, as explained below Eq. S55. In the following, we elaborate on the above:
@@
Here we have split the integral contributions into two parts: the first term is from $\text{Re}\Delta_{K}=\pi$ and exists only when $L$ is odd, while the second term contains all other momentum point contributions. From the arguments below Eq. S55, the first term behaves like:
$\begin{split}\displaystyle\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-(-1)^{L}}}{2L}P^{+-}_{K=\pi+i\frac{\alpha}{L+1}}\text{e}^{i\pi(x_{1}-x_{2})}&\approx-\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-(-1)^{L}}}{2L}\frac{\sqrt{t_{L}/t_{R}-1/t^{2}_{R}}}{2\sqrt{-i\Delta_{K}}}\text{e}^{i\pi(x_{1}-x_{2})}\\ &\xrightarrow[L\rightarrow L_{c}]{\Delta_{K}=K-K_{c},\,\text{Re}\Delta_{K}=0\quad}\quad\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-(-1)^{L}}}{2L}\frac{\sqrt{t_{L}t_{R}-1}}{2t_{R}}\left(\frac{(L+1)(L_{c}+1)}{\alpha(L_{c}-L)}\right)^{1/2}\text{e}^{i\pi(x_{1}-x_{2})}\\ &\xrightarrow{L\approx\sqrt{(L_{c}+1)(L+1)}}\qquad\quad({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-(-1)^{L}})\frac{\sqrt{t_{L}t_{R}-1}}{4\sqrt{\alpha}\cdot t_{R}}(L_{c}-L)^{-1/2}\text{e}^{i\pi(x_{1}-x_{2})}\ .\\ \end{split}$ (S60)
@@
In the above, we have not focused on the spatial $x_{1}-x_{2}$ dependence, since we are primarily concerned about the $L$-scaling behavior. Diagonalizing $\bar{P}$, it is numerically verified [See Fig. S4(a)] that due to our scaling-induced EP, odd and even-sized systems behave qualitatively differently. Near the critical size, the EE (computed from the eigenvalues $p$ of $\bar{P}$ via Eq. S57) for odd $L$ exhibits a divergence which we call an entanglement dip, while in even-$L$ systems, it changes continuously with size.
However, this entanglement dip disappears if the system does not possess such a scaling-dependent GBZ. If we were to instead consider Eq. S46 to be the physical (not effective) model for a non-antagonistic NHSE chain which exhibits a scaling-independent EP point, such that $K=k+i\frac{\alpha}{L_{0}+1}$ with fixed $L_{0}$ (instead of $L$), its EE would no longer show abrupt changes with size, despite still exhibiting odd/even system size effects [See Fig. S4(b)].
Figure S5: Entanglement dip with different entanglement cuts. For different truncated regions $[x_{L},x_{R}]$ implemented by the operator ¯${\bar{\mathcal{R}}}_{[x_{L},x_{R}]}=\sum^{L}_{x\notin[x_{L},x_{R}]}|x\rangle\langle x|\otimes\mathbb{I}$, similar behavior is observed in the entanglement entropy (EE) of (a) our effective 2-component model $H(K)$ Eq. S46 and (b) its parent 4-component coupled chain model $\mathcal{H}$ Eq. S1. For all cases, a prominent entanglement dip occurs, with approximate asymptotic $(L_{c}-L)^{-0.5}$ behavior. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3},\delta=1.6\times 10^{-3}$, corresponding to $L_{c}=42.79$ for (a) and $L^{\prime}_{c}=41.8$ for (b).

### v2

@@ I.2 The new approach: Scale-dependent energies and GBZ
    This approximate expression for the GBZ is dependent on the system size $L$ in the form of the exponent $1/(L+1)$, which gives weak dependence on the inter-chain coupling $\delta$ as well as a constant factor containing the other system parameters. While it is not the only possible approximation, it is relatively compact and agrees well with the numerical GBZ values (Fig. S1b) obtained by substituting the exact OBC spectrum $E_{0}$ of our 4-component physical system into $\text{det}(\mathcal{H}(z)-E_{0}\mathbb{I})=0$, and solving for $z$. Also, it predicts a spectrum $E_{\text{GBZ}}=\pm\sqrt{(t_{L}+1/z_{1})(t_{R}+z_{1})}$ that agrees very well with that numerically-obtained exact OBC spectrum $E_{0}$, as shown in Fig. S1c-f. Regardless of the exact choice of approximation, the $\delta^{1/(L+1)}$ dependence holds universally, and suggests a slow but $\delta$-dependent convergence of the GBZ onto the unit circle in the $L\rightarrow\infty$ limit. In Fig. S1c’-f’, the $|z_{1}|^{x},|z_{2}|^{x}$ spatial decay profiles predicted by the GBZ (brown) near $E_{0}=0$ also exhibit excellent agreement with the upper and lower envelope limits of the numerical eigenstate $\Psi$ closest to $E_{0}=0$ in Fig. S1c-f.
For different values of $\delta$ and system sizes $L$, we assume that the GBZ changes slowly and continuously. In all, our approximate size-dependent GBZ solution for our coupled HN model system is given by
@@
is the transition length in which the energy spectrum or skin depth $-\log|z|$ switches from being size-independent to size-dependent (Fig. S1b). The value of $L^{\prime}$ is determined by setting both expressions for the GBZ to be equal i.e. $|z_{1}|=\sqrt{t_{R}/t_{L}}=\exp\left(-{\alpha}/({L^{\prime}+1})\right)$. The very good agreement between our analytic GBZ solution (Eq. (S44)) and exact numerics is presented in Fig. S1 for the skin depth $-\log|z|$, spectrum $E$ and eigenfunction decay.
@@
Figure S1: Scaling-dependent GBZ, spectrum and eigenfunctions of the 4-component model $\mathcal{H}$ (Eq. (S1)). (a) The two chain 4-component model (Eq. (S1)) incorporates equal but opposite hopping asymmetries $t_{L}$ and $t_{R}$ in both chains I and II. Even a small inter-chain coupling $\delta$ can lead to significant scaling dependence when $t_{L}\neq t_{R}$. (b) Agreement of the scale-dependent skin depth $\log|z|$ obtained via two different ways: numerically, by solving $\det(\mathcal{H}(z)-E_{0})=0$ from the numerical eigenspectrum $E_{0}$ and choosing slowest decaying $z$ solution (blue crosses); analytically, via the scale-dependent generalized Brillouin zone (GBZ) we derived (Eq. (S44)) (red dots) . $L^{\prime}$ (Eq. (S45)) represents the system size at which the skin depth (GBZ) transitions from being scale-independent to scale-dependent. (c-f, c’-f’) Energy spectra at various system sizes $L$ and illustrative eigenfunctions at energies $E_{0}$ (green diamonds in (c-f)) of smallest absolute value. Excellent agreement is observed between the numerical (blue) and analytically-derived GBZ (red) spectra (Eq. (S44)) and eigenfunctions (Eq. (S22), Eq. (S23)). The lower bound for the eigenfunctions (lower red curves) results from the superposition of $z_{1}$ and $z_{2}$ solutions in Eq. (S23), and is generally non-monotonic except at the critical scale $L\approx L_{c}$ (Eq. (LABEL:Eqcritical1)), where it exhibits a perfectly exponential profile since both $z$ solutions coalesce. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$, and $\delta=1.6\times 10^{-3}$ for all panels.
@@
In a nutshell, we have shown that as the system size $L$ increases, the GBZ and spectrum always switches from being size-independent to size-dependent as $L$ crosses $L^{\prime}$. This is given by our approximate GBZ expression Eq. (S44) which nevertheless exhibits excellent agreement with numerical results from exact diagonalization. Our approach was based on neglecting the effect of weak coupling terms in the bulk equations but retaining their influence in enforcing the boundary conditions, which allowed for analytic headway in writing down the energy expression from the characteristic equation of a single chain $\det(H(z)-E)=0$. This is justified because while extremely weak couplings should not significantly affect the bulk, they would still nontrivially interplay with the boundary conditions due to the high sensitivity of non-Hermitian systems to spatial inhomogeneities. To conclude, with nonzero coupling between chains with antagonistic NHSE, the GBZ is found to be no longer given by the usual $|z_{1}|=|z_{2}|$ condition, but instead takes on a $L$-dependent behavior that crucially involves the coupling strength $\delta$.
@@ I.3 Effective size-dependent two-component model
As such, working in our effect size-dependent GBZ $z$, we can accurately study $\mathcal{H}$ (Eq. (S1)) as two uncoupled copies of 2-component chains $H$ i.e. $\mathcal{H}=H(K)\oplus H^{T}(-K)$, where $K=-i\log z$. The effects of the coupling and NHSE are implicit in the imaginary part of $K$. From Eq. (S2),
@@
from Eq. (S44), where the real part $k=\text{Re}K=2m\pi/(L+1)$ with $m=1,2,...,L$, and the imaginary part $\text{Im}K=\displaystyle\frac{\alpha}{L+1}$ is controlled by
@@
from Eqs. (S44) and (S45). From direct diagonalization of $H(K)$, the eigenenergies are given by
@@
which has a similar form as Eq. (S3); but when evaluated on the GBZ (Eq. (S44)), it should reproduce the OBC spectrum of the entire original 4-component model $\mathcal{H}$ (Eq. (S1)). Note that we have used the notation $\mathcal{E}$ to refer to the eigenenergies predicted by the GBZ, which are supposed to closely approximate the true eigenenergies $E$.
@@
with $t_{R},t_{L}>0$ and $\alpha=-\log\left|\delta\times\displaystyle\frac{t_{R}(t_{R}-t_{L})}{t_{L}(t_{R}-1)}\right|$. Here, $L_{c}$ represents the critical length in which an EP appears, and should not be confused with $L^{\prime}=-\alpha/\log\sqrt{t_{R}/t_{L}}-1$ (Eq. (S45)), which is the length above which the GBZ becomes size dependent. Evidently, such a scaling-induced EP exists only if $L_{c}>L^{\prime}$.
@@
Figure S2: The critical system size $L_{c}$ dependence on $\delta$ for the coupled 4-component model (Eq. (S1)) and its effective two-component approximation (Eq. (S46)). The spectral gap of our model closes at an EP at the critical system size $L_{c}$, which is analytically predicted to depend logarithmically on the small inter-chain coupling parameter $\delta$. Excellent agreement is observed between the numerically determined integer $L_{c}$ (black dots) and its analytical prediction (Eq. (LABEL:Eqcritical1)). Since $\delta$ is built into the GBZ for the two-component, single-chain approximation (Eq. (S46)), its excellent numerical agreement is testimony to the validity of our scale-dependent GBZ approach.
@@ I.3.1 Critical behavior near the scaling-induced EP
To prepare for the study of the entanglement properties near the scaling-induced EPs described above, we describe how the Hamiltonian behaves in their neighborhood. We first expand the size-dependent quasi-momentum $K$, both its real and imaginary parts, as $\Delta_{K}=K-K_{c}$ where $K=k+i\alpha/(L+1)$ and $K_{c}=\pi+i\alpha/(L_{c}+1)$ (See Eqs. (S44) and (LABEL:Eqcritical1)), such that $\Delta_{K}=\text{Re}\Delta_{K}+i\text{Im}\Delta_{K}$ is given by
@@
With these, the Hamiltonian (Eq. (S46)) can be expanded near $K=K_{c}$ as
$\begin{split}H(K)&\xrightarrow{\Delta_{K}=K-K_{c}}\begin{pmatrix}0&t_{L}+\text{e}^{-iK_{c}}\text{e}^{-i\Delta_{K}}\\ t_{R}+\text{e}^{iK_{c}}\text{e}^{i\Delta_{K}}&0\end{pmatrix}\\ &\xrightarrow[t_{R}+\exp(iK_{c})=0]{\text{e}^{i\Delta_{K}}\approx 1+i\Delta_{K}}\begin{pmatrix}0&t_{L}-1/t_{R}\\ -it_{R}\Delta_{K}&0\end{pmatrix}\ ,\\ \end{split}$ (S52)
due to $t_{L}-1/t_{R}\neq 0$. This yields the eigenenergies of $H(K_{c}+\Delta_{K})$ to be ±=±$\pm\mathcal{E}=\pm\sqrt{i(1-t_{L}t_{R})\Delta_{K}}$.
@@ I.3.2 Critical properties of the occupied band projector $P_{K}$
[…]
@@
where $\mathcal{E}(K)$ is the eigenvalue of the Hamiltonian $H(K)$ (Eq. (S46)). Note that we have analytically continued $P_{k}\rightarrow P_{K}$ into the GBZ. Ordinarily, away from gap closure points, $P(K)$ projects onto a well-defined occupied band since $\mathcal{E}(K)\neq 0$. But at an EP ($K=K_{c}$)where $\Delta_{K}=0$, $\mathcal{E}(K)=\sqrt{i(1-t_{L}t_{R})\Delta_{K}}$ vanishes, rendering $P_{K}$ singular.
Closest to an EP, where both the real and imaginary parts of $\Delta_{K}$ is very small,
@@
Figure S3: Potentially divergent matrix element $P_{K}^{+-}$ (Eq. (S54)) of our truncated projection operator $P_{K}$ (Eq.(S53)) in even vs. odd $L$ systems. (a) Complex values of the off-diagonal element $P_{K(k)}^{+-}$, which traces a different loop for each system size $L$ as $k$ cycles over a period $[0,2\pi)$ with discrete points $(2m-1)\pi/L$, $m=1,2,...,L$ (drawn as solid dots). The $k=\pi$ point exist only for odd $L$, and causes $P_{K(k)}^{+-}$ to diverge to infinity as $L\rightarrow L_{c}=-\frac{\alpha}{\log t_{R}}-1=42.79$ (Eq. (LABEL:Eqcritical1)), $\alpha=-\log\left|\delta\times\displaystyle\frac{t_{R}(t_{R}-t_{L})}{t_{L}(t_{R}-1)}\right|$. (b, c) The values of $|P_{K(k)}^{+-}|$, Re$(P_{K(k)}^{+-})$ and Im$(P_{K(k)}^{+-})$ as a function of system size $L$, plotted as separate branches for each $k$ value for (b) even and (c) odd $L$. While there exists for even $L$ a $k=\pi-\pi/(L+1)$ branch that departs from the other branches, for odd $L$, the $k=\pi$ branch exhibits a sharp kink for $L=43\approx L_{c}=42.79$. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3},\delta=1.6\times 10^{-3}$.
@@ II Free-fermion entanglement entropy dip due to scaling-induced EPs
where ¯$\bar{\mathcal{R}}_{[x_{L},x_{R}]}$ is the projector onto sites outside of $[x_{L},x_{R}]$ (note that the index $n$ was expressed as indices $(k,\mu)$ in the main text, such as to emphasize the momentum $k$). For two-component models, the above general expression for $P$ reduces to Eq. (S53). The eigenvalues of the projection operators can be interpreted as occupancy probabilities. For $P$, they are limited to either 1 and 0. However, when the system is divided in to the $[x_{L},x_{R}]$ and its complement, the truncated occupied matrix $\bar{P}$ reveals the entanglement between these two regions, resulting in occupancy probabilities (eigenvalues) $p$ away from $1$ and 00. The free-fermion entanglement entropy (EE) $S$ quantifies this entanglement, and is given by
@@
II.1 Entanglement entropy scaling and dip in the effective 2-component model Eq. (S46)
@@
[…]
Here we have split the integral contributions into two parts: the first term is from $\text{Re}\Delta_{K}=\pi$ and exists only when $L$ is odd, while the second term contains all other momentum point contributions. From the arguments below Eq. (S55), the first term behaves like
$\begin{split}\displaystyle\frac{{1-(-1)^{L}}}{2L}P^{+-}_{K=\pi+i\frac{\alpha}{L+1}}\text{e}^{i\pi(x_{1}-x_{2})}&\approx-\frac{{1-(-1)^{L}}}{2L}\frac{\sqrt{t_{L}/t_{R}-1/t^{2}_{R}}}{2\sqrt{-i\Delta_{K}}}\text{e}^{i\pi(x_{1}-x_{2})}\\ &\xrightarrow[L\rightarrow L_{c}]{\Delta_{K}=K-K_{c},\,\text{Re}\Delta_{K}=0\quad}\quad\frac{{1-(-1)^{L}}}{2L}\frac{\sqrt{t_{L}t_{R}-1}}{2t_{R}}\left(\frac{(L+1)(L_{c}+1)}{\alpha(L_{c}-L)}\right)^{1/2}\text{e}^{i\pi(x_{1}-x_{2})}\\ &\xrightarrow{L\approx\sqrt{(L_{c}+1)(L+1)}}\qquad\quad({1-(-1)^{L}})\frac{\sqrt{t_{L}t_{R}-1}}{4\sqrt{\alpha}\cdot t_{R}}(L_{c}-L)^{-1/2}\text{e}^{i\pi(x_{1}-x_{2})}\ .\\ \end{split}$ (S60)
@@
In the above, we have not focused on the spatial $x_{1}-x_{2}$ dependence, since we are primarily concerned about the $L$-scaling behavior. Diagonalizing $\bar{P}$, it is numerically verified (See Fig. S4a) that due to our scaling-induced EP, odd and even-sized systems behave qualitatively differently. Near the critical size, the EE (computed from the eigenvalues $p$ of $\bar{P}$ via Eq. (S57)) for odd $L$ exhibits a divergence which we call an entanglement dip, while in even-$L$ systems, it changes continuously with size.
However, this entanglement dip disappears if the system does not possess such a scaling-dependent GBZ. If we were to instead consider Eq. (S46) to be the physical (not effective) model for a non-antagonistic NHSE chain which exhibits a scaling-independent EP point, such that $K=k+i\frac{\alpha}{L_{0}+1}$ with fixed $L_{0}$ (instead of $L$), its EE would no longer show abrupt changes with size, despite still exhibiting odd/even system size effects (See Fig. S4b).
Figure S5: Entanglement dip with different entanglement cuts. For different truncated regions $[x_{L},x_{R}]$ implemented by the operator ¯${\bar{\mathcal{R}}}_{[x_{L},x_{R}]}=\sum^{L}_{x\notin[x_{L},x_{R}]}|x\rangle\langle x|\otimes\mathbb{I}$, similar behavior is observed in the entanglement entropy (EE) of (a) our effective 2-component model $H(K)$ Eq. (S46) and (b) its parent 4-component coupled chain model $\mathcal{H}$ Eq. (S1). For all cases, a prominent entanglement dip occurs, with approximate asymptotic $(L_{c}-L)^{-0.5}$ behavior. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3},\delta=1.6\times 10^{-3}$, corresponding to $L_{c}=42.79$ for (a) and $L^{\prime}_{c}=41.8$ for (b).